OpenStudy (anonymous):

How to solve 2e^ln x + 3 = 6?

OpenStudy (anonymous):

first off \[e^{\ln(x)}=x\]

OpenStudy (anonymous):

because the log and the exponential functions are inverses it is not clear exactly if you meant \[2e^{\ln(x)=3}=6\] or \[2e^{\ln(x)}+3=6\]

OpenStudy (anonymous):

the answer will be different depending on which it is is it the first one of the second one?

OpenStudy (anonymous):

It would be the second one. @satellite73